OpenStudy (anonymous):

Suppose p varies directly as d, and p = 2 when d = 7. What is the value of d when p = 10?

OpenStudy (anonymous):

@aussy123

OpenStudy (anonymous):

Im copying this off of yahoo. Im not really sure of the question but this may help. When two numbers vary directly to each other, it means they equal each other with a constant mixed in. We'll call that constant "k". d = kp Since d = 7 when p = 2, we can solve for k 7 = 2k k = 2/7 so the equation is d = 2/7 p When we plug in 10 for p, it becomes d = 2/7 * 10 = 20/7 2. Just plug in 2. T(2) = 2 * 8 = 16 3. The absolute value doesn't matter in this one because you're plugging 4 for x, and 4 is a positive number. f(4) = |4| * 7 = 28

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

@brookie5150

OpenStudy (anonymous):

d+35

OpenStudy (anonymous):

oops, d=35.

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

no problem, let me know if you have any others.

OpenStudy (anonymous):

OK